TAB2 needs help in converting general form into slope intercept form. Please show step by step notes. Please help me to determind Ax + By = C
1.General Form Equation is 3x – 2y = 6 into a slope intercept form.

Answer by Alan3354(69443)

TAB2 needs help in converting general form into slope intercept form. Please show step by step notes. Please help me to determind Ax + By = C
-----------------
I don't know what you mean by "TAB2"
-----------------
1.General Form Equation is 3x – 2y = 6 into a slope intercept form.
Putting into slope-intercept means solve for y.
3x – 2y = 6
-2y = -3x + 6
y = (3/2)x - 3 (slope-intercept form)